Two cars start from the same point at the same time and travel in opposite directions. The fastest car travels 12 miles per hour faster than the other. If after 6 hours, the cars are 648 km apart, what is the speed of each car?
Found 2 solutions by vasumathi, amoresroy:
Answer by vasumathi(46)

Let the speed of the slower car = s
then the speed of the faster car = s + 12
so within one hour the distance between them is
d = s+12 + s
d = 2s + 12
In 6 hours the distance would be 6d = 6(2s+12)
so....
6(2s+12 ) = 648
divide both sides by 6
2s+12 = 648/6
2s+12 = 108
subtract 12 on both sides
2s = 108-12
2s = 96
divide both sides by 2
s = 48
so the speed of the slower car = 48 mph
and the speed of the faster car = 48 + 12 = 60 mph

Answer by amoresroy(361)

Two cars start from the same point at the same time and travel in opposite directions. The fastest car travels 12 miles per hour faster than the other. If after 6 hours, the cars are 648 km apart, what is the speed of each car?
Let x = the speed of the slower car
x + 12= the speed of the faster car
6(x + x+12) = 648
12x+72= 648
12x = 576
x = 48
x+12 = 48+12 = 60
The speed of the slower car is 48mph while the speed of the faster car is 60mph.